OpenStudy (anonymous):

Find the slope of the line: 2x-y=1

OpenStudy (anonymous):

please show work

OpenStudy (anonymous):

2

OpenStudy (anonymous):

howd u get that

OpenStudy (anonymous):

looked at equation. asked my potato.

OpenStudy (anonymous):

u sure thats right?

OpenStudy (anonymous):

my potato nodded.

OpenStudy (anonymous):

Rolled a die and got 2. 2 is the answer

OpenStudy (anonymous):

y = mx + b slope-intercept form, the coefficient on x is the slope

OpenStudy (anonymous):

Take 2x - y = 1 subtract 2x from both sides -y = -2x + 1 Divide by -1 y = 2x - 1 yes 2 is the answer

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

fluttering hell. my potato can do maths.

OpenStudy (anonymous):

When a linear equation is written in the form y = mx + b m is the slope. Therefore, the slope of this line is 2

OpenStudy (anonymous):

Have you got it?